OpenStudy (shadowlegendx):

Math SSAT Question The height of a maple tree and a cherry tree have a ratio of 5:2. If the maple tree grew 20cm and 20cm was cut off of the top of the cherry tree, the ratio of their heights would be 3:1. How much taller is the maple tree than the cherry tree. A. 240cm B. 160cm C. 400cm D. 280cm E. 260cm

OpenStudy (shadowlegendx):

@mathslover @Mertsj

OpenStudy (shadowlegendx):

Please explain the answer :)

OpenStudy (anonymous):

when we say "the height of a maple tree and a cherry tree have a ratio of 5:2" we mean that: $$M=\frac52C$$when we read "the maple tree grew 20cm and 20cm was cut off of the top of the cherry tree, the ratio of their heights would be 3:1", that means that the ratio between the new maple tree height \(M+20\) and the new cherry tree height \(C-20\) is 3:1 i.e. $$M+20=\frac31(C-20)$$

OpenStudy (anonymous):

now that we have a system of equations, you can use typical methods -- elimination, substitution, etc. -- to solve the system: $$M+20=3(C-20)\\\frac52C+20=3C-60\\20=\frac12C-60\\80=\frac12C\\C=160$$

OpenStudy (anonymous):

i.e. the cherry tree is 160 cm tall. can you find the height of the maple tree now?

OpenStudy (shadowlegendx):

The maple tree is 400cm and it is 240cm taller than the cherry tree

OpenStudy (anonymous):

very good :-)

OpenStudy (shadowlegendx):

@oldrin.bataku Why would you put the ratio times the cherry tree instead of the maple. What is the rational? Im talking about the M= 5/2C part.

OpenStudy (anonymous):

the rational \(5/2\) corresponds to the ratio \(5:2\); when we say the ratio of A to B is 1/2, for example, we mean \(A=\frac12B\)

OpenStudy (anonymous):

in our case it is stated that the ratio is of the height of a maple tree *to* that of a cherry tree -- note the order of maple, cherry. this is our hint that we write \(M=\frac52C\)

OpenStudy (shadowlegendx):